Find the general integrals of the PDE

(Y+1)p+(x+1)q=z
1
Expert's answer
2021-07-20T12:37:17-0400

"\frac{dx}{y+1}=\frac{dy}{x+1}=\frac{dz}{z}"


"\int (x+1)dx=\int (y+1)dy"


"x^2\/2+x=y^2\/2+y+c_1"


"\frac{dx-dy}{y-x}=\frac{dz}{z}"


"-ln(x-y)=ln(c_2z)"


"c_2=\frac{1}{z(x-y)}"


"F(x^2/2+x-y^2/2-y,\frac{1}{z(x-y)})=0"